You run a construction firm. You have just won a contract to build a government office complex. Building it will require an investment of $9.5 million today and $5.3 million in one year. The government will pay you $20.5 million in one year upon the​ building's completion. Suppose the interest rate is 10.7%.

a. What is the NPV of this​ opportunity?

b. How can your firm turn this NPV into cash​ today?

a. What is the NPV of this​ opportunity?

The NPV of the proposal is? million. (Round to two decimal​ places.)
Answer #1

1.
=-9.5-5.3/1.107+20.5/1.107=4.230803975 million

2.
Borrow 4.23 million now and pay 4.23*1.107=4.68 million 1 year later